# Details Enhancement in Unsigned Distance Field Learning for High-fidelity 3D Surface Reconstruction

### v1

@@ Abstract
While Signed Distance Fields (SDF) are well-established for modeling watertight surfaces, Unsigned Distance Fields (UDF) broaden the scope to include open surfaces and models with complex inner structures. Despite their flexibility, UDFs encounter significant challenges in high-fidelity 3D reconstruction, such as non-differentiability at the zero level set, difficulty in achieving the exact zero value, numerous local minima, vanishing gradients, and oscillating gradient directions near the zero level set. To address these challenges, we propose Details Enhanced UDF (DEUDF) learning that integrates normal alignment and the SIREN network for capturing fine geometric details, adaptively weighted Eikonal constraints to address vanishing gradients near the target surface, unconditioned MLP-based UDF representation to relax non-negativity constraints, and a UDF-tailored method for extracting iso-surface with non-constant iso-values. These strategies collectively stabilize the learning process from unoriented point clouds and enhance the accuracy of UDFs. Our computational results demonstrate that DEUDF outperforms existing UDF learning methods in both accuracy and the quality of reconstructed surfaces. We will make the source code publicly available.
1 Introduction
@@
While signed distance fields (SDF) are favored for their capability to represent watertight surfaces, unsigned distance fields (UDF) provide a means to model both open surfaces and objects with complex inner structures. However, achieving high-quality UDFs that accurately reconstruct 3D surfaces with fine geometric details is challenging for several reasons. Firstly, UDFs struggle to precisely achieve a zero value, making it difficult to identify the exact surface boundaries. Secondly, UDFs are theoretically non-differentiable at the zero level set, resulting in vanishing gradients near the target surface. This issue leads to numerous undesired local minima, complicating the extraction of the zero level set. Thirdly, the gradient directions of UDFs tend to oscillate near the surface, causing the reconstructed surfaces to be fragmented [1].
@@
Due to the inherently low accuracy of learned UDFs, the extracted zero level sets are typically over-smoothed and lack crucial geometric details. Several studies have aimed to enhance the precision of UDF learning. For instance, NDF [2] trains a shape encoder and a decoder from 3D surfaces of various types, including point clouds, meshes and mathematical functions. As a supervised method, its performance heavily relies on the quality and diversity of the training dataset. Unsupervised approaches, such as CAP-UDF [3] and LevelSetUDF [4], offer greater flexibility in handling a wider range of 3D models. Despite advancements in UDF learning techniques, all existing methods still suffer from relatively low accuracy in the learned distance fields compared to SDFs. This limitation significantly diminishes their practical usage in real-world applications.
@@
This paper introduces a new method, called Details Enhanced UDF (DEUDF) learning, aimed at enhancing the accuracy of UDF learning from unoriented point clouds to ensure that learned UDFs can capture the fine geometric details of target surfaces. A key observation is the significant role normal directions play in learning fine details. Although obtaining globally consistent orientations is challenging due to its combinatorial and global optimization nature, acquiring normal directions locally, for instance, through principal component analysis [5], is feasible. Consequently, we constrain the UDF gradients to align with normal directions to enhance detail capture, while disregarding normal orientations.
@@
To overcome the limitation of UDFs not achieving the exact zero value, we relax the strict requirements that UDFs must be non-negative and that the surface must precisely correspond to the zero iso-surfaces. This adaptation enables the use of an unconditioned multilayer perceptron (MLP) meaning an MLP that outputs its value directly without any additional operations to make the output positive. Unlike traditional methods that generate UDFs by taking the absolute value of a learned SDF [4] – prone to inducing oscillating gradients – or by using the $\mathrm{softplus}$ activation function in MLPs to eliminate negative values [6] – leading to vanishing gradients – our relaxation not only addresses the vanishing gradients but also stabilizes the oscillation of gradient directions near the surface.
@@
While SDFs maintain well-behaved gradients with consistent unit length throughout 3D space, UDFs often experience vanishing gradients at the zero level set, diminishing the effectiveness of uniformly applied Eikonal constraints for UDF learning. To address this issue, we propose an adaptively weighted Eikonal constraint, specifically tailored to align with the unique properties of UDFs. Moreover, we incorporate the SIREN network [7] to represent high-frequency details in UDFs, thereby enhancing the encoding capabilities of our model. We consider the local minimum around zero – both positive and negative – as the intended surface and adopt DCUDF [8], an optimization-based iso-surfacing algorithm, to extract the iso-surface with non-constant iso-values.
@@
By integrating normal alignment, unconditioned MLPs with SIREN activation functions, adaptively weighted Eikonal constraints, and UDF-tailored iso-surfacing techniques, DEUDF significantly improves the accuracy of UDF learning. Evaluations on benchmark datasets demonstrate our method outperforms baseline methods in terms of UDF accuracy and quality of reconstructed surfaces.
2 Related work
@@
Surface reconstruction from point clouds has been studied extensively for the last three decades. The field has seen significant evolution, from computational geometry methods [9, 10] to implicit function techniques [5, 11, 12, 13], and more recently to deep learning approaches [14, 2, 4, 15, 16, 17]. Due to space constraints, this section primarily focuses on deep learning-based 3D reconstruction techniques.
@@
Both signed distance fields and occupancy fields effectively represent closed surfaces. An occupancy field defines whether each point in space is inside or outside a given shape. ONet [18] employs a deep neural network classifier to implicitly represent 3D surfaces as a continuous decision boundary, while IF-Net [19] and CONet [20] use encoders to capture shape. Compared to occupancy fields, SDFs provide additional information about the distance of a point form the surface of the object, making them favored for applications that require accurate shape representation, such as reconstruction, shape interpolation and completion. DeepSDF [14] introduces an innovative implicit encoder that defines the boundary of a 3D shape as the zero level set of a learned implicit function. Following this, numerous neural SDF-based works have been developed. For example, DeepLS [21] utilizes a grid structure to store latent codes for local shape features, SIREN [7] introduces a novel activation function for increasing the network’s capability to capture high-frequency signals, and IDF [22] employs displacement maps to enhance the representation of fine details. Additionally, SDFs have been utilized to represent geometric shapes for neural rendering tasks, such as NeuS[23] and VolSDF [24], which leverage SDFs for 3D reconstruction from multi-view images.
@@
To model general non-watertight surface, Chibane et al. [2] introduced neural unsigned distance fields, which predict the unsigned distance from a query point to the nearest surface point. GIFS [25] models the relationship between points rather than between points and surfaces, while NVF [26] learns a vector as an alternative to calculating gradient from UDFs, representing the direction from query points to the target surface. Unlike these methods, which utilize separate neural networks to extract supplementary information that aids UDF learning, CAP-UDF [3] and GeoUDF [15] focus on enhancing the density of the input point clouds via adopting upsampling techniques. Despite these advancements, the challenge of ambiguous gradients near the zero level set remains, due to the non-differentiability of UDFs at this juncture. To address this challenge, LevelSetUDF[4] introduces constraints between non-differentiable zero level set and differentiable non-zero level set, while DUDF [17] adopts a new representation to maintain differentiability at points close to the target surface. Although LevelSetUDF and DUDF tackle the non-differentiable issues, they still struggle to match the quality of reconstruction–particularly for surfaces with fine details-achieved SDF learning methods. Additionally, similar to SDFs, UDFs are also utilized to implicitly represent 3D shapes in neural rendering tasks, such as 3D reconstruction and novel view synthesis from multi-view images [27, 6, 28]. See Table 1 for a qualitative comparison of existing UDF learning methods.
@@
Extracting the zero level set from UDFs is technically non-trivial, as it is rare for the learned UDFs to precisely reach zero values. There are several research efforts aiming at addressing this issue. Gradient-based methods such as CAP-UDF [3], MeshUDF [1] and GeoUDF [15] use both gradient directions and UDF values to detect zero crossings, while optimization-based techniques, such as DCUDF [8], focus on identifying local minima within the input UDFs.
@@
Figure 1: Illustration of UDF learning with various neural representations. (a) Existing neural network architectures often use an absolute value or softplus function to prevent negative distances. (b) In contrast, our method relaxes the non-negative condition and employs an unconditioned MLP with the SIREN activation function for predicting the distances. (c) To show the differences between existing representations and ours, we consider a plane $\pi$ and a line perpendicular to $\pi$. Points along this line, as they pass through the plane, are used to plot the unsigned distance, shown in (d)-(g). The horizontal axis represents a signed distance range from -0.05 to 0.05, while the vertical axis measures the unsigned distance. Ideally, the unsigned distance should exhibit a perfect “V” shape relative to the signed distance, as shown in (d). However, UDFs parameterized by conditioned MLPs can present defects. For example, learning an SDF followed by taking the absolute value results in a “W”-shaped distance field around the zero level set (e). Employing the softplus activation function to eliminate negative values yields learned UDFs with vanishing gradients across a relatively large distance range near the zero value (f). In contrast, our method, which employs unconditioned MLPs, significantly narrows this range of vanishing gradients (g).
3 Method
@@
3.1 Relaxation of non-negative constraints
@@
Observing both vanishing gradients and oscillating gradient directions stem from the strict non-negative constraint on distance values, we propose relaxing the conditions that require UDFs to be non-negative and the surface to coincide precisely with zero iso-surface. We use unconditioned MLPs to represent UDFs and consider the local minimum of the UDF value around zero, which may be either positive or negative, as the point through which the intended surface passes. As illustrated in Figure 1 (e), this relaxation results in a distance function with a significantly narrower bandwidth compared to using the softplus activation function, thereby providing a high-quality approximation to the ground truth distance filed, which exhibits a V-shaped profile.
@@
In NeuralUDF [27], a similar loss term in the form $\exp(-100|f|)$ was used. It is important to note that our loss term does not include the absolute value. This subtle difference significantly impacts the behavior of the learned distance field $f$. With the absolute value, their loss encourages $|f|$ being a large positive value, which consequently reduces the occurrence of points with zero distance values. This reduction minimizes the presence of small disconnected components in the reconstructed surfaces [22, 27]. Therefore, the $\exp$ in their loss functions acts as a regularizer to smooth the learned distance fields. As mentioned above, the use of the absolute value $|f|$ in the loss function can lead to undesired side effects, such as a W-shaped profile in the learned UDFs, which may consequently result in watertight models. In sharp contrast, our loss term, which omits the absolute value, serves as a soft non-negative constraint. This encourages $f$ to remain positive as much as possible, thus differentiating it from an SDF, and enabling $f$ to mimic a true UDF. Even though DCUDF [8] used an unconditional MLP to represent UDF, it needs the ground truth UDF for supervision. Otherwise, it cannot learn the UDF.
3.2 Normal alignment
Normal directions are critical for enhancing surface details in the reconstruction process. Let $\mathcal{N}=\{\mathbf{n}_{i}\}_{i=1}^{n}$ represent the set of unit normals for the point set $\mathcal{P}$. Following [5], we apply principal component analysis to each point $\mathbf{p}_{i}$ to determine its normal direction $\mathbf{n}_{i}$. Since UDF gradients typically vanish on the surface, it is impractical to directly constrain the gradients of $\mathcal{P}$.
To address this issue, we generate a sample point set $\mathcal{Q}=\{\mathbf{q}_{i}\}_{i=1}^{n}$ in each training epoch, where each point $\mathbf{q}_{i}$ is strategically displaced from the surface. Specifically, $\mathbf{q}_{i}=\mathbf{p}_{i}+\lambda_{i}\mathbf{n}_{i}$, with the displacement $\lambda_{i}$ randomly chosen from the ranges $[-0.003,0]$ and $[0,0.003]$, respectively. This ensures that $\mathcal{Q}$ contains samples on both sides of the surface, enabling a balanced evaluation of regions close to the geometric structure of interest. We then impose constraints on the UDF gradient directions at points in $\mathcal{Q}$, calculated using the following normal alignment loss term:
$\mathcal{L}_{\mathrm{normal}}=\sum_{\mathbf{q}_{i}\in\mathcal{Q},\mathbf{n}_{i}\in\mathcal{N}}\left(1-\frac{\nabla f(\mathbf{q}_{i})\cdot\mathbf{n}_{i}}{\|\nabla f(\mathbf{q}_{i})\|_{2}\cdot\|\mathbf{n}_{i}\|_{2}}\mathrm{sign}(\lambda_{i})\right).$ (3)
3.3 Adaptively weighted Eikonal constraints
@@
where the weight function, $\delta(\cdot)$, is designed to reduce the contribution from points close to the target surface. A U-shaped function with controllable bandwidth serves this purpose effectively. In our implementation, we employ the attenuation function used in IDF [22] as our weight as $\delta(d)=\left(1+(\frac{\xi}{d})^{4}\right)^{-1}$, where $\xi$ represents the threshold beyond which the influence of the Eikonal constraint begins to diminish significantly. In our experiments, we initially set $\xi$ to 0.01 and gradually decrease $\xi$ to 0.002 over the course of the learning process, following the learning rate. This adjustment is made to enhance the attenuation effect. We evaluate the Eikonal loss $\mathcal{L}_{\mathrm{eikonal}}$ for points in the set $\mathcal{Q}$, which serves as a proxy of the target geometry, as well as for randomly sampled points throughout the entire domain $\Omega$.
3.4 UDF learning
For the network architecture, we employ a 5-layer SIREN network [7], which consists of 256 units per layer. The network utilizes a sinusoidal activation function $sin(\omega x)$, with a frequency parameter $\omega=60$ in our implementation, to effectively encode high-frequency details. Our network takes spatial coordinates $(x,y,z)$ as inputs and outputs the predicted unsigned distance.
The training process aims to minimize the following loss function:
@@ 3.4 UDF learning
where $\lambda_{i}$s are weights assigned to balance the contributions from the four loss terms. We empirically set $\lambda_{1}=400$, $\lambda_{2}=50$, $\lambda_{3}=40$ and $\lambda_{4}=10$ in our implementation. We train the neural network using the Adam [29] optimizer, starting with a learning rate of 0.00005. The learning rate decays to zero following a cosine annealing[30] schedule.
3.5 Isosurface extraction
After obtaining the UDFs, we proceed to extract the iso-surface from the learned UDFs. Due to the relaxation of the non-negative constraints, the target geometry does not align precisely with the zero level set. Instead, we identify the target surface as the one passing through the local minima (which can be either positive or negative) of the UDFs near the zero values.
One possible method for extracting the target geometry from UDFs involves explicitly using the UDF gradient, such as MeshUDF [1] and GeoUDF [15], both of which are variants of the standard Marching Cubes algorithm. On each cube edge, if the gradient directions of the UDF at the two endpoints are opposite and their UDF values are below a specified threshold, a zero crossing is marked on that edge. However, this approach is not suitable for our purpose because our learned UDFs do not ensure the positivity of the distance value, and the local minima may exceed the specified UDF threshold. If this occurs, the iso-surfacing method would exclude these cubes, leading to the creation of extracted surfaces with undesired holes.
(a) DUDF
(b) NSH
(c) LevelSetUDF
(d) Ours
(e) GT
Figure 2: Visual comparison with two recent UDF learning approaches, DUDF [17] and LevelSetUDF [4], and one recent SDF learning method, NSH [31], on surfaces with fine geometric details. Our method yields visually pleasing results, reconstructing significantly more details than the other methods.
To tackle this challenge, we adopt the optimization-based method DCUDF [8], which initiates by extracting a double cover using the Marching Cubes algorithm at a small positive iso-value on the UDF. Subsequently, it shrinks the double cover to the local minimum of the UDF. This method does not require the UDF to be strictly positive nor does it depend on a threshold to select candidate cubes. As a result, it effectively identifies the local minima, yielding a high-quality triangle mesh that accurately represents the target surface.
4 Experiments
4.1 Datasets
We evaluate our method using three datasets: ShapeNet-Cars [32] with 108 models (We select all models named starting with “1”), the Stanford 3D Scene Dataset [33] with 5 models and the Stanford 3D Scan Repository  [^1] https://graphics.stanford.edu/data/3Dscanrep/ with 8 models. For each shape, we randomly sample 300K points as input. After learning the UDFs, we employ DCUDF with a resolution of $512^{3}$ to extract the target surface. To evaluate the accuracy, We use Chamfer distance (CD) and F-score as quantitative measures. For F-score, we set the thresholds to 0.01 and 0.005. Following previous methods [3, 4], we randomly sample 100K points from both the reconstructed surfaces and the ground truth meshes for computing CD and F-score. We report our results on an NVIDIA Tesla V100 GPU with 32GB memory (about 5GB used for a UDF learning). It takes about 30 minutes to learn a UDF.
4.2 Results & comparisons
We compare our method with three state-of-the-art UDF learning methods: LevelSetUDF [4], CAP-UDF [3] and DUDF[17]. Since we adopt DCUDF [8] for surface extraction, we also test DCUDF for the three baselines to ensure fairness of comparisons. For CAP-UDF [3] and LevelSetUDF [4], we observe that DCUDF could produce better results than their original implementations in terms of Chamfer distances and visual effects. But for DUDF, the results of DCUDF are not as good as the original results. Therefore, to report the best results of the baseline methods, we choose to adopt DCUDF for extracting the zero level set from the UDF outputs from both CAP-UDF and LevelSetUDF. While DUDF uses their original results for comparisons. For completeness, we also report the original results of CAP-UDF and LevelSetUDF in the supplementary material for visual and quantitative comparisons.
Additionally, we assess our approach against IDF [22] and NSH [31], two state-of-the-art SDF learning methods, on the watertight surfaces with fine geometric details from the Stanford 3D Scan Repository. See Table 2 and Figure 2.
3D objects with fine geometric details
3D objects with complex inner structures
We further explore our method for representing 3D objects and scenes with complex inner structures. We evaluate our method on Stanford 3D Scene and ShapeNet-Cars datasets. As shown in Table 2 and Figure 3, our method is more stable on complex structures and performs optimally in the representation of open boundaries.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
(a) DUDF
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
(b) LevelSetUDF
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
(c) Ours
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
(d) GT
Figure 3: Visual comparison with DUDF and LevelSetUDF on an indoor scene from the Stanford 3D Scene dataset featuring noise and an imperfect scan, and two car models from the ShapeNet-Cars dataset showcasing complex inner structures. The open boundaries generated by our method are the best, such as, the building windows, inner structures of the vehicle and car exhaust vent.
4.3 Ablation studies
We conduct ablation studies to demonstrate the effectiveness of each component within our method. The results are shown in Figure 4 and Table 3.
Table 2: Quantitative results on the Stanford 3D Scene Dataset, the Stanford 3D Scans (watertight) and ShapeNet-Cars. Chamfer distances are measured in the unit of ×$\times 10^{-3}$. We randomly sampled 100K points from each ground truth mesh and then computed both CDs and F-scores as reference performance metrics. Therefore, the closer the actual performance metrics are to these references, the higher the quality of the results produced by the method. For CAP-UDF and LevelSetUDF, following their official surface extraction implementations, we removed artifacts that are far from the ground truth point cloud after extracting the mesh by DCUDF.

### v2

@@ Abstract
While Signed Distance Fields (SDF) are well-established for modeling watertight surfaces, Unsigned Distance Fields (UDF) broaden the scope to include open surfaces and models with complex inner structures. Despite their flexibility, UDFs encounter significant challenges in high-fidelity 3D reconstruction, such as non-differentiability at the zero level set, difficulty in achieving the exact zero value, numerous local minima, vanishing gradients, and oscillating gradient directions near the zero level set. To address these challenges, we propose Details Enhanced UDF (DEUDF) learning that integrates normal alignment and the SIREN network for capturing fine geometric details, adaptively weighted Eikonal constraints to address vanishing gradients near the target surface, unconditioned MLP-based UDF representation to relax non-negativity constraints, and DCUDF for extracting the local minimal average distance surface. These strategies collectively stabilize the learning process from unoriented point clouds and enhance the accuracy of UDFs. Our computational results demonstrate that DEUDF outperforms existing UDF learning methods in both accuracy and the quality of reconstructed surfaces.
Introduction
@@
While signed distance fields (SDF) are favored for their capability to represent watertight surfaces, unsigned distance fields (UDFs) provide a means to model both open surfaces and objects with complex inner structures. However, achieving high-quality UDFs that accurately reconstruct 3D surfaces with fine geometric details is challenging for several reasons. Firstly, UDFs struggle to precisely achieve a zero value, making it difficult to generate open boundaries. Secondly, UDFs are theoretically non-differentiable at the zero level set, resulting in vanishing gradients near the target surface. This issue leads to numerous undesired local minima, complicating the extraction of the zero level set. Thirdly, the gradient directions of UDFs tend to oscillate near the surface, causing the reconstructed surfaces to be fragmented (Guillard, Stella, and Fua 2022).
@@
Due to the inherently low accuracy of learned UDFs, the extracted zero level sets are typically over-smoothed and lack crucial geometric details. Several studies have aimed to enhance the precision of UDF learning. For instance, NDF (Chibane, Mir, and Pons-Moll 2020) trains a shape encoder and a decoder from 3D surfaces of various types, including point clouds, meshes and mathematical functions. As a supervised method, its performance heavily relies on the quality and diversity of the training dataset. Unsupervised approaches, such as CAP-UDF (Zhou et al. 2022) and LevelSetUDF (Zhou et al. 2023), offer greater flexibility in handling a wider range of 3D models. Despite advancements in UDF learning techniques, all existing methods still suffer from relatively low accuracy in the learned distance fields compared to SDFs. This limitation significantly diminishes their practical usage in real-world applications.
@@
This paper introduces a new method, called Details Enhanced UDF (DEUDF) learning, aimed at enhancing the accuracy of UDF learning from unoriented point clouds to ensure that learned UDFs can capture the fine geometric details of target surfaces. A key observation is the significant role normal directions play in learning fine details. Although obtaining globally consistent orientations (Xu et al. 2023) is challenging due to its combinatorial and global optimization nature, acquiring normal directions locally, for instance, through principal component analysis (Hoppe et al. 1992), is feasible. Consequently, we constrain the UDF gradients to align with normal directions to enhance detail capture, while disregarding normal orientations.
@@
Existing methods strictly constrain UDFs to non-negative values, so the UDFs are hard to achieve exact zero values. Meanwhile, such strict constraint has brought some other problems as illustrated in Figure 1. To overcome this limitation, we relax the strict requirements that UDFs must be non-negative and that the surface must precisely correspond to the zero iso-surfaces. Though the distances are not strictly non-negative, we still call it UDF. This adaptation enables the use of an unconditioned multilayer perceptron (MLP), meaning an MLP that outputs its value directly without additional operations to make the output positive. Unlike traditional methods that generate UDFs by taking the absolute value of a learned SDF (Zhou et al. 2023) – prone to inducing oscillating gradients – or by using the $\mathrm{softplus}$ activation function in MLPs to eliminate negative values (Liu et al. 2023) – leading to vanishing gradients – our relaxation not only addresses the vanishing gradients but also stabilizes the oscillation of gradient directions near the surface.
@@
While SDFs maintain well-behaved gradients with consistent unit length throughout 3D space, UDFs often experience vanishing gradients at the zero level set, diminishing the effectiveness of uniformly applied Eikonal constraints for UDF learning. To address this issue, we propose an adaptively weighted Eikonal constraint, specifically tailored to align with the unique properties of UDFs. Moreover, we incorporate the SIREN network (Sitzmann et al. 2020) to represent high-frequency details in UDFs, thereby enhancing the encoding capabilities of our model. We consider the intended surface should be around zero and the distance values—both positive and negative—should be as small as possible. To achieve this, we adopt DCUDF (Hou et al. 2023), an optimization-based surface extraction algorithm.
@@
By integrating normal alignment, unconditioned MLPs with SIREN activation functions, adaptively weighted Eikonal constraints, and UDF-tailored surface extraction techniques, DEUDF significantly improves the accuracy of UDF learning. Evaluations on benchmark datasets demonstrate our method outperforms baseline methods in terms of UDF accuracy and quality of reconstructed surfaces.
Related work
@@
Surface reconstruction from point clouds has been studied extensively for the last three decades. The field has seen significant evolution, from computational geometry methods (Amenta and Bern 1998; Dey and Goswami 2003) to implicit function techniques (Hoppe et al. 1992; Ohtake et al. 2003; Kazhdan, Bolitho, and Hoppe 2006; Kazhdan and Hoppe 2013; Hou et al. 2022; Liu et al. 2024), and more recently to deep learning approaches (Park et al. 2019; Chibane, Mir, and Pons-Moll 2020; Zhou et al. 2023; Ren et al. 2023; Wang et al. 2023a; Fainstein, Siless, and Iarussi 2024). Due to space constraints, this section primarily focuses on deep learning-based 3D reconstruction techniques.
@@
Both signed distance fields and occupancy fields effectively represent closed surfaces. An occupancy field defines whether each point in space is inside or outside a given shape. ONet (Mescheder et al. 2019) employs a deep neural network classifier to implicitly represent 3D surfaces as a continuous decision boundary, while IF-Net (Chibane, Alldieck, and Pons-Moll 2020) and CONet (Peng et al. 2020) use encoders to capture shape. Compared to occupancy fields, SDFs provide additional information about the distance of a point form the surface of the object, making them favored for applications that require accurate shape representation, such as reconstruction, shape interpolation and completion. DeepSDF (Park et al. 2019) introduces an innovative implicit encoder that defines the boundary of a 3D shape as the zero level set of a learned implicit function. Following this, numerous neural SDF-based works have been developed. For example, DeepLS (Chabra et al. 2020) utilizes a grid structure to store latent codes for local shape features, SIREN (Sitzmann et al. 2020) introduces a novel activation function for increasing the network’s capability to capture high-frequency signals, and IDF (Wang, Rahmann, and Sorkine-Hornung 2022) employs displacement maps to enhance the representation of fine details. Additionally, SDFs have been utilized to represent geometric shapes for neural rendering tasks, such as NeuS (Wang et al. 2021) and VolSDF (Yariv et al. 2021), which leverage SDFs for 3D reconstruction from multi-view images.
@@
To model general non-watertight surface, Chibane et al. (Chibane, Mir, and Pons-Moll 2020) introduced neural UDFs, which predict the unsigned distance from a query point to the nearest surface point. GIFS (Ye et al. 2022) models the relationship between points rather than between points and surfaces, while NVF (Yang et al. 2023) learns a vector as an alternative to calculating gradient from UDFs, representing the direction from query points to the target surface. Unlike these methods, which utilize separate neural networks to extract supplementary information that aids UDF learning, CAP-UDF (Zhou et al. 2022) and GeoUDF (Ren et al. 2023) focus on enhancing the density of the input point clouds via adopting upsampling techniques. Despite these advancements, the challenge of ambiguous gradients near the zero level set remains. To address the non-differentiability issue of UDFs at zero, LevelSetUDF (Zhou et al. 2023) projects the properties of the non-zero level set to the zero level set to learn a continuous and smooth UDF, while DUDF (Fainstein, Siless, and Iarussi 2024) adopts a new representation to maintain differentiability at points close to the target surface. Although LevelSetUDF and DUDF tackle the non-differentiable issues, they still struggle to match the quality of reconstruction–particularly for surfaces with fine details-achieved SDF learning methods. Additionally, similar to SDFs, UDFs are also used to implicitly represent 3D shapes in neural rendering tasks, such as 3D reconstruction from multi-view images (e.g., NeuralUDF (Long et al. 2023), NeUDF (Liu et al. 2023), 2S-UDF (Deng et al. 2024)) and 3D generation (Yu et al. 2025; Zhou et al. 2024). See Table 1 for a qualitative comparison of existing UDF learning methods.
@@
Extracting the zero level set from UDFs is technically non-trivial, as it is rare for the learned UDFs to precisely reach zero values. There are several research efforts aiming at addressing this issue. Gradient-based methods such as CAP-UDF (Zhou et al. 2022), MeshUDF (Guillard, Stella, and Fua 2022) and GeoUDF (Ren et al. 2023) use both gradient directions and UDF values to detect zero crossings, while optimization-based techniques, such as DCUDF (Hou et al. 2023), optimize the surface locally to minimize the average distance value on the surface.
@@
Figure 1: Illustration of UDF learning with various neural representations. (a) Existing neural network architectures often use an absolute value or softplus function to prevent negative distances. (b) In contrast, our method relaxes the non-negative condition and employs an unconditioned MLP with the SIREN activation function for predicting the distances. (c) To show the differences between existing representations and ours, we consider a plane $\pi$ and a line perpendicular to $\pi$. We plot the unsigned distance to $\pi$ on this line learned by our method, but with different MLP output layers. (d) The horizontal axis represents a signed distance range from -0.05 to 0.05, while the vertical axis measures the learned unsigned distance. Ideally, the unsigned distance should exhibit a perfect “V” shape relative to the signed distance. (e) However, UDFs parameterized by conditioned MLPs can present defects. For example, learning a negative value followed by taking the absolute value results in a “W”-shaped distance field around the zero level set. (f) Employing the softplus activation function to eliminate negative values yields learned UDFs with vanishing gradients across a relatively large distance range near the zero value. (g) In contrast, our method, which employs unconditioned MLPs, significantly narrows this range of vanishing gradients.
Method
@@
Relaxation of non-negative constraints
@@
Observing both vanishing gradients and oscillating gradient directions stem from the strict non-negative constraint on distance values, we propose relaxing the conditions that require UDFs to be non-negative and the surface to coincide precisely with zero iso-surface. We use unconditioned MLPs to represent UDFs and consider the local minimal distance surface of the UDF value around zero, which may be either positive or negative, as the intended surface. As illustrated in Figure 1 (e), this relaxation results in a distance function with a significantly narrower bandwidth compared to using the softplus activation function, thereby providing a high-quality approximation to the ground truth distance filed, which exhibits a V-shaped profile.
@@
In NeuralUDF (Long et al. 2023), a similar loss term in the form $\exp(-100|f|)$ was used. It is important to note that our loss term does not include the absolute value. This subtle difference significantly impacts the behavior of the learned distance field $f$. With the absolute value, their loss encourages $|f|$ being a large positive value, which consequently reduces the occurrence of points with zero distance values. This reduction minimizes the presence of small disconnected components in the reconstructed surfaces (Wang, Rahmann, and Sorkine-Hornung 2022; Long et al. 2023). Therefore, the $\exp$ in their loss functions acts as a regularizer to smooth the learned distance fields. As mentioned above, the use of the absolute value $|f|$ in the loss function can lead to undesired side effects, such as a W-shaped profile in the learned UDFs, which may consequently result in watertight models. In sharp contrast, our loss term, which omits the absolute value, serves as a soft non-negative constraint. This encourages $f$ to remain positive as much as possible, thus differentiating it from an SDF, and enabling $f$ to mimic a true UDF. Even though DCUDF (Hou et al. 2023) used an unconditional MLP to represent UDF, it needs the ground truth UDF to supervise and keep the distances almost positive. While, DEUDF is unsupervised learning.
Normal alignment
Normal directions are critical for enhancing surface details in the reconstruction process. Let $\mathcal{N}=\{\mathbf{n}_{i}\}_{i=1}^{n}$ represent the set of unit normals for the point set $\mathcal{P}$. Following (Hoppe et al. 1992), we apply principal component analysis to each point $\mathbf{p}_{i}$ to determine its normal direction $\mathbf{n}_{i}$. Since UDF gradients typically vanish on the surface, it is impractical to directly constrain the gradients of $\mathcal{P}$.
To address this issue, we generate a set of sample point pairs $\mathcal{Q}=\{(\mathbf{q}^{1}_{i},\mathbf{q}^{2}_{i})\}_{i=1}^{n}$ in each training epoch, where each point $\mathbf{q}_{i}$ is strategically displaced from the surface. Specifically, $\mathbf{q}^{1}_{i}=\mathbf{p}_{i}+\lambda_{i}\mathbf{n}_{i}$ and $\mathbf{q}^{2}_{i}=\mathbf{p}_{i}-\lambda_{i}\mathbf{n}_{i}$, with the displacement $\lambda_{i}$ being randomly chosen in the range $(0,0.003]$. This ensures that $\mathcal{Q}$ contains samples on both sides of the surface, enabling a balanced evaluation on both sides of the geometric structure of interest. We then impose constraints on the UDF gradient directions at points in $\mathcal{Q}$, using the following normal alignment loss term:
$\mathcal{L}_{\mathrm{normal}}=\sum_{\stackrel{{\scriptstyle\mathbf{n}_{i}\in\mathcal{N},}}{{k\in\{1,2\}}}}\left(1+(-1)^{k}\frac{\nabla f(\mathbf{q}^{k}_{i})\cdot\mathbf{n}_{i}}{\|\nabla f(\mathbf{q}^{k}_{i})\|_{2}\cdot\|\mathbf{n}_{i}\|_{2}}\right).$ (3)
Adaptively weighted Eikonal constraints
@@
where the weight function, $\delta(\cdot)$, is designed to reduce the contribution from points close to the target surface. A U-shaped function with controllable bandwidth serves this purpose effectively. In our implementation, we employ the attenuation function used in IDF (Wang, Rahmann, and Sorkine-Hornung 2022) as our weight as $\delta(d)=\left(1+(\frac{\xi}{d})^{4}\right)^{-1}$, where $\xi$ represents the threshold beyond which the influence of the Eikonal constraint begins to diminish significantly. In our experiments, we initially set $\xi$ to 0.01 and gradually decrease $\xi$ to 0.002 over the course of the learning process, following the learning rate. This adjustment is made to enhance the attenuation effect. We evaluate the Eikonal loss $\mathcal{L}_{\mathrm{eikonal}}$ for points in the set $\mathcal{Q}$, which serves as a proxy of the target geometry, as well as for randomly sampled points throughout the entire domain $\Omega$.
UDF learning
For the network architecture, we employ a 5-layer SIREN network (Sitzmann et al. 2020), which consists of 256 units per layer. The network utilizes a sinusoidal activation function $\sin(\omega x)$, with a default frequency [^1] For noisy point clouds, using a lower frequency enhances resilience against noise. $\omega=60$ for clean point clouds in our implementation, effectively encoding fine geometric details. Our network takes spatial coordinates $(x,y,z)$ as inputs and outputs the predicted unsigned distance. The training process aims to minimize the following loss function:
@@
where $\lambda_{i}$s are weights assigned to balance the contributions from the four loss terms. We empirically set $\lambda_{1}=400$, $\lambda_{2}=50$, $\lambda_{3}=40$ and $\lambda_{4}=10$ in our implementation. We train the neural network using the Adam (Kingma and Ba 2015) optimizer, starting with a learning rate of $5\times 10^{-5}$. The learning rate decays to zero following a cosine annealing schedule (Loshchilov and Hutter 2017).
@@
Method Distance Mean Median $F1^{0.005}$ $F1^{0.0025}$ Mean Median $F1^{0.005}$ $F1^{0.0025}$ Mean Median $F1^{0.005}$ $F1^{0.0025}$